# LaV-CoT: Language-Aware Visual CoT with Multi-Aspect Reward Optimization for Real-World Multilingual VQA

### v3

@@ 3.3. Model Training
We adopt Supervised Fine-Tuning (SFT) to equip the model with robust multilingual and multimodal reasoning abilities. To further refine output quality, we introduce a composite reward function, $R_{\mathrm{Multi\_Aspect}}$, which aggregates multiple rule-based criteria into a single supervisory signal. Instead of relying on an explicit value function or critic, candidate outputs are comparatively evaluated within the same group, and their raw rewards are transformed into relative advantage scores. This pairwise normalization stabilizes optimization and promotes the preference for higher-quality responses.
The multi-aspect reward $R_{\mathrm{Multi\_Aspect}}$ consists of four complementary components: three novel rewards designed to capture different aspects of multilingual visual-textual reasoning, and a default format reward that ensures the generated output adheres to the expected structural conventions.
@@
$R_{\mathrm{Count}}=1-\frac{|N_{Ts}-\hat{N}_{Ts}+N_{Obj}-\hat{N}_{Obj}|}{N_{Ts}+N_{Obj}}$
@@
The GRPO objective evaluates how well the current policy generates outputs compared to the previous policy. Probabilities under the current and previous policies are compared for importance sampling, while a clipping mechanism and KL-divergence penalty ensure stable updates and prevent excessive deviation from a reference policy. Each candidate’s advantage is computed relative to others in the same group using multi-aspect rewards, promoting the selection of higher-quality outputs without a separate value function. By integrating rewards for language consistency, text segmentation, object count, answer correctness, and output format, GRPO provides rich, fine-grained supervision for end-to-end multilingual multimodal reasoning.
Figure 3. Comparison between Qwen2.5-VL-7B and LaV-CoT. As illustrated, Qwen2.5-VL-7B demonstrates a step-by-step reasoning process; however, it fails to perform the reasoning in the target Arabic language and produces an incorrect final deduction. In contrast, LaV-CoT effectively follows its reasoning pipeline to produce the correct final answer.:
@@ 4. Experiment
In this section, we first provide an overview of the experimental framework, including implementation details, evaluation benchmarks, and the MLLMs used for comparison. We then present a comprehensive comparison of LaV-CoT with state-of-the-art approaches on multilingual benchmarks, followed by an evaluation across a variety of multimodal tasks. Finally, we conclude with ablation studies and qualitative visualization of multilingual cases, highlighting the exceptional reasoning capabilities of our model.
@@ 4.1. Experiment Setting
Implementation Details.
During the Supervised Fine-Tuning (SFT) stage, the model was trained for 1–3 epochs with a global batch size of 256 on 8 NVIDIA A100 (40GB) GPUs. The image resolution size is set to be 896 * 896. The learning rate was fixed at $2\times 10^{-4}$ and kept constant without decay, using the fused AdamW optimizer. Mixed-precision training with bfloat16 (bf16) and TensorFloat-32 (tf32) was enabled to improve computational efficiency while maintaining numerical stability. Gradient clipping with a maximum norm of 0.3 and a warm-up ratio of 0.03 were applied. Parameter-efficient fine-tuning was implemented via LoRA(Hu et al., 2021), targeting the transformer modules q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj, with rank 32, alpha 64, and a dropout rate of 0.05. In the subsequent GRPO training stage, all reward components were uniformly scaled by 0.25 so that the total reward summed to 1, i.e.,
@@
Evaluation Benchmark.
For our experiments, MMMB(Sun et al., 2024), Multilingual MMBench(Sun et al., 2024), and MTVQA(Tang et al., 2024) serve as the primary evaluation benchmarks. Both MMMB and Multilingual MMBench focus on assessing the general multilingual visual question answering (VQA) capabilities of models, covering a wide range of question types and multimodal reasoning skills. While MTVQA is designed to specifically measure a model’s ability to align textual and visual information in text-centric VQA scenarios, with an emphasis on accurately recognizing and interpreting embedded text within images. Together, these datasets provide a complementary evaluation setting that captures both broad reasoning abilities and fine-grained multimodal text understanding. For evaluation, we adopt VLMEvalKit (Duan et al., 2024) with consistent configuration settings, thereby ensuring a fair and consistent comparison.
Comparison Models.
We adopt Qwen2.5-VL-3B as our backbone model and train it under the proposed LaV-CoT framework. For comprehensive evaluation, we compare against a diverse set of multilingual multimodal models (MLLMs), covering different model scales and accessibility levels. Specifically, we include: (1) open-source models of comparable size, such as Qwen2-VL-2B(Wang et al., 2024a), InternVL3-2B(Zhu et al., 2025), InternVL3.5-2B(Wang et al., 2025) and Qwen2.5-VL-3B (Bai et al., 2025) as the baseline; (2) larger open-source models with roughly double the parameters, including Monkey(Li et al., 2024b), LLaVA-OneVision(Li et al., 2024c), PARROT(Sun et al., 2024), DeepSeek-VL-7B(Lu et al., 2024), Qwen2-VL-7B(Wang et al., 2024a), Qwen2.5-VL-7B(Bai et al., 2025), InternVL3-8B(Zhu et al., 2025) and InternVL3.5-8B(Wang et al., 2025); (3) much larger proprietary frontier models, such as GPT-4o-0513(OpenAI, 2023) and Gemini-2.5-flash(DeepMind, 2023). In addition, we also report results of BlueLM-V-3B(Xiong et al., 2025), which, although closed-source, demonstrates strong multilingual and multimodal capabilities.
@@
LaV-CoT (SFT) 83.2 81.1 80.6 80.5 77.3 82.2 85.7 84.4 84.3 83.9 85.2 83.1 19.8 33.7 38.1 33.6 20.7 37.1 10.4 22.1 37.7 64.7
LaV-CoT (SFT + GRPO) 86.0 84.8 83.0 82.7 80.3 83.6 89.0 89.4 85.9 88.3 86.8 87.7 23.2 35.3 42.8 35.9 23.1 38.2 11.5 23.8 38.8 67.5
Table 1. Multilingual Vision-Language Reasoning Model Results. Results are presented for MMMB, Multilingual MMBench, and MTVQA across multiple languages.
@@ 5.1. Quantitative Analysis
Table 1 presents the multilingual evaluation results of a wide range of vision-language models across three benchmarks: MMMB, Multilingual MMBench, and MTVQA. Several important observations can be drawn.
@@
Overall and comparative performance. Large proprietary models such as GPT-4o and Gemini-2.5-flash achieve strong performance across most benchmarks, with overall accuracy of 66.6 and 67.4, respectively, establishing a strong upper bound for multilingual multimodal reasoning. Among open-source models, Qwen2.5-VL-7B, InternVL3.5-8B, and our proposed LaV-CoT variants deliver competitive performance, achieving overall scores of 64.4, 64.9 and 67.5, respectively. Notably, LaV-CoT (SFT + GRPO) surpasses all open-source baselines: on MMMB and Multilingual MMBench it performs strongly across most languages except for slightly lower scores on PT and RU, while on MTVQA it achieves particularly high accuracy for AR and KO.
@@
Medium-sized models such as Monkey, DeepSeek-VL-7B, and GLM-4v-9B remain below 50 overall, highlighting the challenges of multilingual multimodal reasoning in low-resource settings. Smaller-scale models like Qwen2.5-VL-3B, InternVL3-2B and InternVL3.5-2B achieve modest improvements above 50, whereas larger open-source systems such as PARROT, LLaVA-OneVision, and Qwen2-VL-7B deliver substantially stronger results. Our LaV-CoT (SFT) model outperforms strong baselines including Qwen2.5-VL-7B and InternVL3.5-8B, and with reinforcement learning fine-tuning (SFT+GRPO), it achieves state-of-the-art performance, approaching or slightly exceeding proprietary models.
Impact of GRPO. Importantly, incorporating GRPO training further improves LaV-CoT. The LaV-CoT (SFT + GRPO) variant achieves an overall score of 67.5, outperforming all open-source models and even surpassing the proprietary Gemini-2.5-flash and GPT-4o on specific multilingual settings (e.g., Arabic, Turkish, Korean, etc.). This highlights the effectiveness of reinforcement learning with preference optimization in enhancing cross-lingual reasoning.
@@
Benchmark-specific insights. On MMMB and Multilingual MMBench, LaV-CoT achieves robust improvements across multiple languages, particularly in Arabic and Turkish, where traditional baselines exhibit significant drops. On MTVQA, our model also delivers substantial gains in low-resource languages such as Arabic, Korean, and Vietnamese, showcasing stronger cross-lingual visual reasoning ability.
Overall, these results confirm that LaV-CoT, especially when enhanced with GRPO, achieves state-of-the-art performance among open-source multilingual multimodal models and narrows the gap with proprietary systems.
Figure 4. Training Reward Curves:
5.2. GRPO Training Reward Analysis
@@
Figure 4 illustrates the smoothed reward curves during GRPO training, highlighting the evolution of four key reward components: Language Reward, Count Reward, Edit Distance Reward, and Format Reward. The Format Reward exhibits rapid initial improvement, ascending from approximately 0.125 to 0.25 within the first 850 steps before stabilizing, indicating the base model has desent instruction following capability and shows early convergence in format adherence. In contrast, the Language Reward experiences a brief decline from 0.13 to 0.11 by step 800, followed by a steady ascent to 0.25 around step 6500, with subsequent fluctuations reflecting ongoing refinement in linguistic quality. The Count Reward demonstrates a gradual rise from 0.06, including a plateau between steps 3000 and 4000, reaching stability at 0.24 post-step 6540, suggesting progressive optimization of quantitative accuracy. The Edit Distance Reward initially decreases from 0.13 to 0.10 by step 2600, then recovers to 0.22 by step 7500 and converges around 0.23, underscoring challenges in minimizing textual discrepancies early in training. Curves are smoothed using a moving average with a window size of 10, and shaded areas represent ±1 standard deviation, reflecting the variability in reward estimates. Overall, these dynamics demonstrate GRPO’s effectiveness in balancing multi-objective rewards, achieving high-fidelity convergence by the end of training.
Model $\text{OCR}_{avg}$ $\text{SR}_{avg}$ $\text{IR}_{avg}$ $\text{OL}_{avg}$ $\text{SIU}_{avg}$
Qwen2.5-VL-3B (Baseline) 83.7 68.5 74.4 72.2 75.0
Qwen2.5-VL-3B (Direct Train) 85.7 71.5 76.4 74.4 77.0
Qwen2.5-VL-3B (Text CoT) 87.7 73.5 78.4 76.1 79.0
LaV-CoT (SFT) 88.6 75.2 80.5 81.0 81.5
Table 2. Ablation study with different training methods.
Figure 5. Abliation Study on LaV-CoT GRPO Training.:
6. Abliataion Study
@@ 6.1. Impact of Language-Aware Visual CoT on Performance
Table 2 presents an ablation study comparing different training strategies for Qwen2.5-VL-3B. The baseline model is trained without CoT, while ”Direct Train” denotes standard supervised fine-tuning, and ”Text CoT” incorporates textual Chain-of-Thought reasoning. LaV-CoT (SFT) further integrates language-aware visual CoT, resulting in consistent improvements across all evaluation metrics. To evaluate the effectiveness of each method, we examine several representative tasks from Multilingual MMBench: Spatial Relationship (SR), Identity Reasoning (IR), Object Localization (OL), and Structured Image-Text Understanding (SIU). Notably, LaV-CoT achieves the highest scores across OCR, SR, IR, OL, and SIU, demonstrating the substantial benefits of incorporating structured multilingual visual reasoning into the model.
6.2. Rewards Ratio Ablation Study
To investigate the impact of different reward configurations on model performance, we conduct an ablation study by varying the reward ratios $(\alpha,\beta,\gamma,\delta)$ in the GRPO framework. Figure 5 illustrates the performance across nine representative languages on MTVQA. The baseline configuration $(0,0,0,0)$, which does not incorporate any reward signals, achieves the lowest performance, with an average accuracy of 28.90. Incorporating format-only rewards $(0,0,0,1)$ yields a minimal improvement over the baseline. Introducing partial structural rewards $(0,0,0.5,0.5)$ results in a moderate gain, increasing the average accuracy to 29.36. A more balanced reward setting $(0,0.33,0.33,0.33)$ further improves performance, achieving an average accuracy of 30.11. Finally, the fully balanced configuration $(0.25,0.25,0.25,0.25)$ delivers the best overall results, with an average accuracy of 31.13, demonstrating the effectiveness of jointly optimizing multiple reward dimensions. These findings highlight that carefully balancing the GRPO rewards is crucial for enhancing multilingual reasoning robustness.

### v4

@@ 3.2. Model Training
3.2.1. Two-stage Training Paradigm
In the first stage, we adopt SFT to equip the model with basic multilingual and multimodal reasoning abilities, followed by GRPO-based RL post-training in the second stage to further enhance its performance. However, standard GRPO rewards are typically coarse-grained and binary (e.g., assigning 1 for a correct answer and 0 otherwise), failing to capture the fine-grained reasoning quality required for multilingual TEC-VQA. Moreover, conventional GRPO provides no explicit supervision over the reasoning process itself, such as maintaining language consistency or ensuring coherent step-by-step reasoning quality. To further refine output quality, we introduce a composite reward function, $R_{\mathrm{Multi\_Aspect}}$, which aggregates multiple rule-based criteria into a single supervisory signal.
3.2.2. Multi-aspect Reward
The multi-aspect reward $R_{\mathrm{Multi\_Aspect}}$ consists of four complementary components: three novel rewards designed to capture different aspects of multilingual visual-textual reasoning, and a format reward that ensures the generated output adheres to the expected structural conventions.
@@ 3.2.2. Multi-aspect Reward
$R_{\mathrm{Count}}=\begin{cases}1-\dfrac{\left|(N_{Ts}-\hat{N}_{Ts})+(N_{Obj}-\hat{N}_{Obj})\right|}{N_{Ts}+N_{Obj}},&\text{if }N_{Ts}+N_{Obj}>0,\\ 1,&\text{otherwise}\end{cases}$
@@
The GRPO objective evaluates how well the current policy generates outputs compared to the previous policy. Probabilities under the current and previous policies are compared for importance sampling, while a clipping mechanism and KL-divergence penalty ensure stable updates and prevent excessive deviation from a reference policy. By integrating multi-aspect rewards covering language consistency, segmentation, object count, correctness, and format, our refined GRPO provides fine-grained supervision for multilingual multimodal reasoning.
Figure 3. Comparison between Qwen2.5-VL-7B and LaV-CoT. As illustrated, Qwen2.5-VL-7B demonstrates a step-by-step reasoning process; however, it fails to perform the reasoning in the target Arabic language and produces an incorrect final deduction. In contrast, LaV-CoT effectively follows its reasoning pipeline to produce the correct final answer.
demo case
@@ 4. Experiment
In this section, we provide an overview of the experimental framework, including implementation details, evaluation benchmarks, and baseline models used for comparison.
@@ 4.1. Experiment Setting
4.1.1. Implementation Details.
During the Supervised Fine-Tuning (SFT) stage, the model was trained for 1–3 epochs with a global batch size of 256 on 8 NVIDIA A100 (40GB) GPUs. The image resolution size is set to be 896 * 896 (Li et al., 2024b). The learning rate was fixed at $2\times 10^{-4}$ and kept constant without decay, using the fused AdamW optimizer. Mixed-precision training with bfloat16 (bf16) and TensorFloat-32 (tf32) was enabled to improve computational efficiency while maintaining numerical stability. Gradient clipping with a maximum norm of 0.3 and a warm-up ratio of 0.03 were applied. Parameter-efficient fine-tuning was implemented via LoRA(Hu et al., 2021), targeting the transformer modules q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj, with rank 32, alpha 64, and a dropout rate of 0.05. In the subsequent GRPO training stage, all reward components were uniformly scaled by 0.25 so that the total reward summed to 1, i.e.,
@@
4.1.2. Evaluation Benchmark.
For our experiments, MMMB (Sun et al., 2025), Multilingual MMBench (Sun et al., 2025), and MTVQA (Tang et al., 2025) serve as the primary evaluation benchmarks. MTVQA is designed to specifically measure a model’s ability to align textual and visual information in multilingual text-centric VQA scenarios. To verify LaV-CoT’s broader applicability across multilingual VQA scenarios, we additionally evaluate on MMMB and Multilingual MMBench, which target general multilingual multimodal reasoning and a wide variety of question types. Together, these datasets form a complementary evaluation suite that assesses both broad multilingual reasoning capabilities and multilingual multimodal text understanding. For evaluation, we adopt VLMEvalKit (Duan et al., 2025) with consistent configuration settings, thereby ensuring a fair and consistent comparison.
4.1.3. Comparison Models.
We adopt Qwen2.5-VL-3B as our backbone model and train it under the LaV-CoT framework. For a comprehensive evaluation, we compare it against a diverse set of methods, categorized by model type. General models include open-source models of comparable size, such as Qwen2-VL-2B (Wang et al., 2024a), InternVL3-2B (Zhu et al., 2025), InternVL3.5-2B (Wang et al., 2025) and Qwen2.5-VL-3B (Bai et al., 2025), as well as larger open-source models with more than twice the number of parameters, including DeepSeek-VL-7B (Lu et al., 2024), GLM-4v-9B (GLM et al., 2024), Qwen2-VL-7B (Wang et al., 2024a), Qwen2.5-VL-7B (Bai et al., 2025), InternVL3-8B (Zhu et al., 2025) and InternVL3.5-8B (Wang et al., 2025). We also include much larger proprietary frontier models, such as GPT-4o-2024-05-13 (Hurst et al., 2024) and Gemini-2.5-flash (DeepMind, 2023). In addition, we compare with expert models that are either specialized for text-rich visual reasoning and multilingual VQA, which include BlueLM-V-3B (Xiong et al., 2025) Monkey-9.8B (Li et al., 2024b), LLaVA-OneVision-7B (Li et al., 2024a), PARROT-7B (Sun et al., 2025), and Eagle2-9B (Li et al., 2025).
@@
Expert Models
BlueLM-V-3B 17.3 39.5 44.7 32.2 23.5 34.0 9.2 20.3 22.9
Monkey-9.8B 66.0 58.1 46.3 38.8 37.6 48.5 58.0 53.5 49.5 31.0 31.3 45.1 1.1 16.4 21.2 21.7 4.3 5.4 5.3 6.1 8.4 35.0
LLaVA-OneVision-7B 79.0 78.2 75.9 73.3 67.7 76.3 76.7 75.3 73.4 70.4 64.8 73.1 5.0 21.0 22.3 26.1 6.2 7.3 6.0 3.0 13.6 53.7
PARROT-7B 80.1 80.0 79.6 76.5 75.0 79.9 78.0 77.1 76.7 75.9 74.0 77.7 14.3 30.1 34.6 36.7 16.3 20.2 9.5 9.9 27.8 59.7
Eagle2-9B 84.4 83.0 81.8 81.2 75.8 83.0 83.8 83.7 81.5 78.0 75.1 81.1 17.0 29.1 34.8 37.0 17.6 33.3 10.2 13.4 32.2 62.8
Ours
LaV-CoT-3B (SFT) 83.2 81.1 80.6 80.5 77.3 82.2 85.7 84.4 84.3 83.9 85.2 83.1 19.8 33.7 38.1 33.6 20.7 37.1 10.4 22.1 37.7 64.7
LaV-CoT-3B (SFT + GRPO) 86.0 84.8 83.0 82.7 80.3 83.6 89.0 89.4 85.9 88.3 86.8 87.7 23.2 35.3 42.8 35.9 23.1 38.2 11.5 23.8 38.8 67.5
Table 2. Performance of multilingual vision-language models on MMMB, Multilingual MMBench, and MTVQA across multiple languages. For each language and the overall average accuracy, the best results are highlighted in bold and underlined.
@@ 5.1. Quantitative Analysis
Table 2 presents the multilingual evaluation results of various methods based on general and expert vision-language models across three benchmarks: MMMB, Multilingual MMBench, and MTVQA. Several key observations can be made from both overall comparisons and benchmark-specific analyses.
5.1.1. Overall and comparative performance.
@@
Large proprietary models such as GPT-4o and Gemini-2.5-flash achieve strong performance across most benchmarks, with overall accuracy of 66.6% and 67.4%, respectively, establishing a strong upper bound for multilingual multimodal reasoning. Among open-source models, Qwen2.5-VL-7B, InternVL3.5-8B, and our proposed LaV-CoT variants deliver competitive performance, achieving overall scores of 64.4%, 64.9% and 67.5%, respectively. Notably, LaV-CoT (SFT + GRPO) outperforms all baselines on average across MMMB, Multilingual MMBench, and MTVQA.
@@
Medium-sized models such as Monkey-9.8B, DeepSeek-VL-7B, and GLM-4v-9B remain below 50% overall, highlighting the challenges of multilingual multimodal reasoning. Smaller-scale models like Qwen2.5-VL-3B, InternVL3-2B and InternVL3.5-2B achieve modest improvements above 50%, whereas larger open-source systems such as PARROT-7B, LLaVA-OneVision-7B, Qwen2-VL-7B and Eagle2-9B deliver substantially stronger results. Our LaV-CoT (SFT) model outperforms strong baselines, including Qwen2.5-VL-7B and InternVL3.5-8B, and with further reinforcement tuning (SFT + GRPO), it achieves state-of-the-art performance, approaching and slightly exceeding proprietary models.
Figure 4. LaV-CoT GRPO Training Reward Curves.
rewards
5.1.2. Benchmark-specific insights.
@@
On MMMB and Multilingual MMBench, LaV‑CoT achieves robust improvements across most languages, with slightly lower scores on Portuguese and Russian. On MTVQA, our model also delivers substantial gains in low-resource languages such as Arabic and Korean.
Overall, a 3B model trained under the LaV-CoT framework, with multi-aspect reward design and multi-stage reasoning, achieves state-of-the-art performance among open-source multilingual multimodal models and surpasses several proprietary models. We further evaluate additional standard document understanding benchmarks, including DocVQA (Mathew et al., 2021), ChartQA (Masry et al., 2022), and OCRBenchv2 (Fu et al., 2025), with detailed results in Appendix B.
5.2. GRPO Training Reward Curves Analysis
@@
Figure 4 illustrates the smoothed reward curves during GRPO training, highlighting the evolution of four key reward components: Language Reward, Count Reward, Edit Distance Reward, and Format Reward. The Format Reward exhibits rapid initial improvement, ascending from approximately 0.125 to 0.25 within the first 850 steps before stabilizing, indicating the base model has decent instruction following capability and shows early convergence in format adherence. In contrast, the Language Reward experiences a brief decline from 0.13 to 0.11 by step 800, followed by a steady ascent to 0.25 around step 6500, with subsequent fluctuations reflecting ongoing refinement in linguistic quality. The Count Reward demonstrates a gradual rise from 0.06, including a plateau between steps 3000 and 4000, reaching stability at 0.24 post-step 6540, suggesting progressive optimization of quantitative accuracy. The Edit Distance Reward initially decreases from 0.13 to 0.12 by step 2600, then recovers to 0.22 by step 7500 and converges around 0.23, underscoring challenges in minimizing textual discrepancies early in training. Curves are smoothed using a moving average with a window size of 10, and shaded areas represent ±1 standard deviation, reflecting the variability in reward estimates. Overall, these dynamics demonstrate GRPO’s effectiveness in balancing multi-objective rewards, achieving high-fidelity convergence by the end of training.
6. Ablation Study
@@ 6.1. Impact of Language-Aware Visual CoT on Performance
Table 3 presents an ablation study comparing different training strategies with Qwen2.5-VL-3B as the baseline. ”Direct Train” denotes standard supervised fine-tuning on task-specific data without any intermediate reasoning supervision, and ”SFT w/ Text CoT” incorporates textual Chain-of-Thought reasoning. ”SFT w/ LaV-CoT” further integrates language-aware visual CoT, resulting in consistent improvements across representative tasks. On average, fine-tuning with Language-Aware Visual CoT achieves a relative improvement of $\sim$6.6% over the baseline across all tasks, highlighting the effectiveness of structured multilingual visual reasoning.
Table 3. Ablation study of different fine-tuning methods across various tasks on Multilingual MMBench. Here OCR represents Optical Character Recognition, SR represents Spatial Relationship, IR represents Identity Reasoning, OL represents Object Localization, and SIU represents Structured Image-Text Understanding.
Model OCR SR IR OL SIU Avg
Qwen2.5-VL-3B (Baseline) 83.7 68.5 74.4 72.2 75.0 74.8
Qwen2.5-VL-3B (Direct Train) 85.7 71.5 76.4 74.4 77.0 77.0
Qwen2.5-VL-3B (SFT w/ Text CoT) 87.7 73.5 78.4 76.1 79.0 78.9
Qwen2.5-VL-3B (SFT w/ LaV-CoT) 88.6 75.2 80.5 81.0 81.5 81.4
6.2. Comparison of Multi-Aspect Reward Configurations
To investigate the impact of different reward configurations on model performance, we conduct an ablation study by varying the reward ratios $(\alpha,\beta,\gamma,\delta)$ in the GRPO framework. Figure 5 illustrates the performance across nine representative languages on MTVQA. Specifically, we compare four configurations: training with Format Reward $(0,0,0,1)$, which incorporates only the format reward and achieves accuracy of 28.9%; training with Answer & Format Rewards $(0,0,0.5,0.5)$, which covers both answer and format supervision and yields a moderate gain with an average accuracy of 29.36%; training with Count & Answer & Format Rewards $(0,0.33,0.33,0.33)$, which adopts a more balanced setting across count, answer, and format rewards, further improving performance to 30.11%; and training with all four Rewards (LaV-CoT) $(0.25,0.25,0.25,0.25)$, which jointly optimizes all reward dimensions and delivers the best overall results with an average accuracy of 31.13%. These findings highlight that balancing multi-aspect rewards is crucial for robust multilingual multimodal reasoning.